# Can ChatGPT Rival Neural Machine Translation? A Comparative Study

### v1

@@ Abstract
Inspired by the increasing interest in leveraging large language models for translation, this paper evaluates the capabilities of large language models (LLMs) represented by ChatGPT in comparison to the mainstream neural machine translation (NMT) engines in translating Chinese diplomatic texts into English. Specifically, we examine the translation quality of ChatGPT and NMT engines as measured by four automated metrics and human evaluation based on an error-typology and six analytic rubrics. Our findings show that automated metrics yield similar results for ChatGPT under different prompts and NMT systems, while human annotators tend to assign noticeably higher scores to ChatGPT when it is provided an example or contextual information about the translation task. Pairwise correlation between automated metrics and dimensions of human evaluation produces weak and non-significant results, suggesting the divergence between the two methods of translation quality assessment. These findings provide valuable insights into the potential of ChatGPT as a capable machine translator, and the influence of prompt engineering on its performance.
Can ChatGPT Rival Neural Machine Translation? A Comparative Study
Zhaokun ${}^{\dagger}$, Ziyin ${}^{\ddagger}$thanks: daenerystargaryen@sjtu.edu.cn ${}^{\dagger}$ School of Foreign Languages, Shanghai Jiao Tong University ${}^{\ddagger}$ Department of Computer Science and Engineering, Shanghai Jiao Tong University
@@ 1 Introduction
Against the backdrop of translation automation and the advances in translation technology, neural machine translation (hereafter NMT) has been widely investigated under both academic and professional scenarios (Gaspari et al., 2015). Scholars have discussed its applications across different domains, and found it to fare well in both non-literary and literary texts, demonstrating satisfying quality and even creativity to some degree (Hu and Li, 2023).
In recent years, the emergence of large language models (LLMs) has further revolutionized the relationship between translation technologies and translators, and has opened up new possibilities for interactive translation using customized prompts, making proficiency in using AI tools an integral part of digital literacy (Godwin-Jones, 2022). Among them, ChatGPT, an AI-powered LLM designed and trained by OpenAI1, has demonstrated remarkable capabilities in a wide range of tasks, including translation among others (Hendy et al., 2023; Lai et al., 2023; Hendrycks et al., 2021; Srivastava et al., 2023; Zhao et al., 2023; Zhang et al., 2023). Recent studies have shown that ChatGPT is in par with or surpasses mainstream NMT engines in translating news, user comments, conversation, social media posts, and scientific texts (Jiao et al., 2023; Hendy et al., 2023; Wang et al., 2023; Peng et al., 2023; Lu et al., 2023; Chen et al., 2023; He et al., 2023; Kocmi and Federmann, 2023; Jiang et al., 2023).
The earlier explorations focusing on translation quality assessment (TQA) of machine translation have demonstrated the promising role of ChatGPT in generating quality translations. However, most of these studies are conducted on huge translation corpora, and the scholars mostly rely on automated metrics and scarcely resort to human evaluation, which only serves as a complement to automated metrics.
@@
As NMT systems and LLMs keep making improvements, these measures may fail to fully capture the quality of machine translation. In particular, we have limited understanding of what aspects or dimensions other than content overlap can really define high-quality machine translations and distinguish them from those low-quality ones. There is a lack of metrics that are capable of measuring translation quality as humans do, who pay attention to dimensions such as intended audience, cultural sensitivity, adherence to translation norms, textual coherence, and practicality. Also, previous research is often conducted on extensive, publicly available datasets that are crawled from the internet and encompass diverse domains [^1] In this work, the terms “domain” and “’register” are used interchangeably. The former is preferred by the AI community. The later is preferred by linguistics literature., genres, and text types. There seems to be a scarcity of high-quality, domain-specific parallel corpora that would enable more focused and specialized quality assessment in machine translation.
Thus, we focus on a specific domain, and construct our major research questions as follows:
@@
    How well do NMT systems and ChatGPT perform under automated assessment?
@@
    How well do NMT systems and ChatGPT perform under human evaluation?
@@
    To what degree do automated metrics align with human evaluation?
Essentially, RQ1 and RQ2 aim to provide a well-rounded assessment of translations from NMT systems and ChatGPT using both automated metrics and human evaluation. Based on RQ1 and RQ2, RQ3 taps into the inter-correlation between automated metrics and human scores. For RQ1 and RQ2, we find that ChatGPT excels under human evaluation and semantic-aware automatic evaluation (see Figure 1 and Table 2). For RQ3, we find that the two methods may exhibit divergences when approaching machine-translated texts, contrary to Lu and Han (2023)’s findings that automated metrics can show moderate to strong correlations with human-assigned scores in assessing interpreting outputs, possibly due to the inherent differences between interpreting and translation.
@@ 2.1 Improving the Translations of LLM via Prompt Engineering
Large Language Models (LLMs) represented by OpenAI’s ChatGPT have advanced remarkably in recent years, showcasing emergent abilities such as in-context learning and Chain-of-Thought reasoning (Brown et al., 2020; Chowdhery et al., 2022; Wei et al., 2022a, b; Wang et al., 2022), and several studies have explored the influence of prompting strategies on the translation performance of LLMs (Jiao et al., 2023; Hendy et al., 2023; Wang et al., 2023; Peng et al., 2023; Lu et al., 2023; Chen et al., 2023; He et al., 2023). It has been found that with carefully written prompts, LLMs’ translation quality can gain noticeable improvement, though such better performance is not seen in every scenario (Kocmi and Federmann, 2023; Lu et al., 2023). For instance, to inject cultural awareness into LLMs such as ChatGPT, Yao et al. (2023) propose “cultural knowledge prompting”. Likewise, Mu et al. (2023) use translation memories as part of the prompts to enhance the translation performance of LLMs.
@@ 3.1 Corpus
0-shot
@@
1-shot
Here is an example for you:
@@
Context
Please translate the following Chinese sentence {source} into English. It comes from a press conference in which a Chinese official interact with foreign reporters. The Chinese sentence is in the domain of politics. You should only output the translation.
@@ 3.3.3 Human Evaluation Based on Analytic Rubric Scoring
Model
BLEU
Chrf
BERTScore
COMET
ChatGPT 0-shot
23.82
55.83
96.03
84.19
ChatGPT 1-shot
25.08
55.83
96.14
84.64
ChatGPT with context
23.98
56.19
96.38
85.30
Google Translate
25.28
55.86
96.12
83.02
MS Translate
29.16
59.17
96.15
84.41
DeepL
26.39
57.39
96.18
83.92
@@ A.3 Error Typology, Severity Levels and Penalty Levels
Non-translation 100 Deduct the penalty points for non-translation errors
Critical errors 25 Deduct the penalty points for critical errors
Major errors 10 Deduct the penalty points for major errors
Minor errors 1 Deduct the penalty points for minor errors
Neutral errors 0 Deduct the penalty points for neutral errors
@@ Appendix D Average Correlation Coefficients between Human Evaluation and Automated Metrics
Table 14: Correlation coefficients between human evaluation and automated metrics. $*{}^{*}$: $p<0.005$, significant in all six systems.

### v2

# Convergences and Divergences between Automatic Assessment and Human Evaluation: Insights from Comparing ChatGPT-Generated Translation and Neural Machine Translation

@@ Abstract
Large language models have demonstrated parallel and even superior translation performance compared to neural machine translation (NMT) systems. However, existing comparative studies between them mainly rely on automated metrics, raising questions into the feasibility of these metrics and their alignment with human judgment. The present study investigates the convergences and divergences between automated metrics and human evaluation in assessing the quality of machine translation from ChatGPT and three NMT systems. To perform automatic assessment, four automated metrics are employed, while human evaluation incorporates the DQF-MQM error typology and six rubrics. Notably, automatic assessment and human evaluation converge in measuring formal fidelity (e.g., error rates), but diverge when evaluating semantic and pragmatic fidelity, with automated metrics failing to capture the improvement of ChatGPT’s translation brought by prompt engineering. These results underscore the indispensable role of human judgment in evaluating the performance of advanced translation tools at the current stage.
Convergences and Divergences between Automatic Assessment and Human Evaluation: Insights from Comparing ChatGPT-Generated Translation and Neural Machine Translation
Zhaokun Jiang, Ziyin Zhangthanks: daenerystargaryen@sjtu.edu.cn School of Foreign Languages, Shanghai Jiao Tong University Department of Computer Science and Engineering, Shanghai Jiao Tong University
@@ 1 Introduction
Translation quality assessment (TQA) for machine translation gains increasing prominence along with the development of AI-based translation technologies and the increasing popularity of machine translation. Researchers have dedicated extensive theoretical discussions to elucidating the evolving construct of translation quality (Drugan, 2015), and empirical investigations have led to the development of diverse assessment methods to keep abread with the latest machine translation tools (Eyckmans et al., 2016; Chung, 2020; Lu and Han, 2023; Bojar et al., 2016; Popović, 2018; Bentivogli et al., 2018). Yet, it leaves the exploration on the validity of automatic assessment less touched.
Along with the technological turn, neural machine translation (NMT) has witnessed remarkable advancements characterized by reduced grammatical errors and enhanced contextual comprehension (Gaspari et al., 2015; Cho et al., 2014). NMT has been applied across non-literary and literary texts, showcasing satisfactory quality and even a certain degree of creativity (Hu and Li, 2023). The advent of large language models, represented by ChatGPT (Brown et al., 2020; OpenAI, 2023; Chowdhery et al., 2023; Zhang et al., 2023), has further accelerated the progress of translation automation by enabling interactive and customized experiences (Godwin-Jones, 2022). When used for translation tasks, ChatGPT exhibits greater stylistic adaptability and contextual sensitivity in comparison to NMT systems (OpenAI, 2023). This versatility is achieved through the use of written prompts, which are textual instructions to elicit desired outputs or responses from LLMs. Previous research has demonstrated that ChatGPT’s translation performance can vary depending on the prompts employed (Hendy et al., 2023; Jiao et al., 2023; He et al., 2023).
@@
In the context of large-scale and domain-specific MT tasks, automatic assessment has become the prevailing method for TQA due to its efficiency and standardization (Hendy et al., 2023). To evaluate translation quality, a range of automated metrics have been developed. However, it is important to acknowledge that automatic assessment may over-emphasize quantifiable aspects of translation quality while downplaying other subtle yet vital dimensions associated with high-quality translation (O’Brien, 2012; Pym and Torres-Simón, 2021). Specifically, most automated metrics mainly focus on accuracy, measuring the exact matching of n-grams (e.g., BLEU and chfF), or the extent to which the translation accurately reflects the semantic meaning of the reference (e.g., BERTScore and COMET). It is important to consider whether there are any essential dimensions that have been overlooked by automatic assessment and should be integrated into the existing automated methods.
As the counterpart of automatic assessment, human evaluation is widely regarded as the “gold standard” or the “ground truth”, since human raters, especially professional translators, take factors such as cultural appropriateness, style, and norms into consideration. Human evaluation practices encompass a range of methods, including error analysis (Popović, 2018) as well as emerging approaches like rubrics-referenced scoring (Han and Lu, 2023), item-based assessment (Eyckmans et al., 2009), and comparative judgment (Bojar et al., 2016). However, in practice, human evaluation is rarely employed because it is time-consuming, cognitively-taxing, and labor-intensive (Han and Lu, 2023), requiring scoring design, rater recruitment, rater training, and data collection.
Hence, researchers have advocated for the integration of both approaches to overcome the limitations of human evaluation and leverage the strengths of automatic assessment. In pursuit of this objective, Callison-Burch et al. (2006) and Chung (2020) conducted exploratory studies to examine the correlation between automated metrics and human scores. However, their investigations revealed relatively low and statistically insignificant correlation coefficients (below 0.3). These findings suggest that the automated scores are often unreliable and exhibit substantial variation across different contexts. Such inconsistencies may arise from the divergent constructs of translation quality. In other words, human raters may conceptualize translation quality differently from what can be quantified by automated metrics, and there may be significant dimensions within human evaluation that are not captured by automatic scores (Chung, 2020).
Understanding the disparity in how humans and automated systems perceive the construct of translation quality is the crucial first step to leverage their respective advantages. Ideally, given the esteemed status of human evaluation as the gold standard, it is essential to identify the distinctive dimensions inherent in human assessment and incorporate them into existing automated metrics.
The present study thus aims to address this issue by offering preliminary insights and potential solutions through empirical explorations. We conducted automatic assessment and human evaluation on translation outputs generated by three advanced NMT systems and ChatGPT. Special attention was paid to discern the dimensions that distinguish human-assigned scores from automatic scores Our investigation was conducted on a customized corpus consisting of 6,878 Spokesperson’s Remarks. These diplomatic texts possess linguistic complexity and contextual intricacies, making them an ideal testing ground to evaluate how different machine translators handle the nuances and subtleties of language (Jiang et al., 2023). By concentrating on this specific register, we aim to identify dimensions of translation quality that extend beyond accuracy, encompassing dimensions such as cultural awareness and textual coherence. Our ultimate objective is to make automatic assessment more human-like and fine-grained, and to develop a TQA method that exhibits enhanced validity, reliability, and applicability. This endeavor is particularly crucial in light of the emerging challenges posed by the continual advancements in large language models and NMT systems.
We constructed our major research questions (RQs) as follows:
@@
    What do automatic assessment and human evaluation inform us about the translation quality of ChatGPT and NMT?
@@
    To what degree do automated metrics align with human evaluation?
@@
    What are the unique insights from humans in evaluating the performance of advanced translation tools?
@@ 2.1 Improving the Translations of LLM via Prompt Engineering
Large Language Models (LLMs) represented by OpenAI’s ChatGPT have advanced remarkably in recent years, showcasing emergent abilities such as in-context learning and Chain-of-Thought reasoning (Brown et al., 2020; Chowdhery et al., 2023; Wei et al., 2022a, b; Wang et al., 2022), and several studies have explored the influence of prompting strategies on the translation performance of LLMs (Jiao et al., 2023; Hendy et al., 2023; Wang et al., 2023; Peng et al., 2023; Lu et al., 2023; Chen et al., 2023; He et al., 2023). It has been found that with carefully written prompts, LLMs’ translation quality can gain noticeable improvement, though such better performance is not seen in every scenario (Kocmi and Federmann, 2023; Lu et al., 2023). For instance, to inject cultural awareness into LLMs such as ChatGPT, Yao et al. (2023) propose “cultural knowledge prompting”. Likewise, Mu et al. (2023) use translation memories as part of the prompts to enhance the translation performance of LLMs.
@@ 3.1 Corpus
0-shot Please translate the following Chinese sentence {source} into English. You should only output the translation.
@@
1-shot Here is an example for you:
@@
Context Please translate the following Chinese sentence {source} into English. It comes from a press conference in which a Chinese official interact with foreign reporters. The Chinese sentence is in the domain of politics. You should only output the translation.
@@ 3.3.3 Human Evaluation Based on Analytic Rubric Scoring
Model BLEU Chrf BERTScore COMET
ChatGPT 0-shot 23.82 55.83 96.03 84.19
ChatGPT 1-shot 25.08 55.83 96.14 84.64
ChatGPT with context 23.98 56.19 96.38 85.30
Google Translate 25.28 55.86 96.12 83.02
MS Translate 29.16 59.17 96.15 84.41
DeepL 26.39 57.39 96.18 83.92
@@ A.3 Error Typology, Severity Levels and Penalty Levels
Non-translation 100 Deduct the penalty points for non-translation errors Critical errors 25 Deduct the penalty points for critical errors Major errors 10 Deduct the penalty points for major errors Minor errors 1 Deduct the penalty points for minor errors Neutral errors 0 Deduct the penalty points for neutral errors
@@ Appendix D Average Correlation Coefficients between Human Evaluation and Automated Metrics
Table 14: Correlation coefficients between human evaluation and automated metrics. : $p<0.005$, significant in all six systems.